# Estimating the growth rate of a birth and death process using data from a small sample

### v1

@@ Abstract
The problem of estimating the growth rate of a birth and death processes based on the coalescence times of a sample of $n$ individuals has been considered by several authors ([14, 16, 10, 7]). This problem has applications, for example, to cancer research, when one is interested in determining the growth rate of a clone.
Recently, [7] proposed an analytical method for estimating the growth rate using the theory of coalescent point processes, which has comparable accuracy to more computationally intensive methods when the sample size $n$ is large. We use a similar approach to obtain an estimate of the growth rate that is not based on the assumption that $n$ is large.
We demonstrate, through simulations using the R-package cloneRate, that our estimator for the growth rate has a smaller mean squared error than previous estimates when $n$ is small.
@@ 1 Introduction
Consider a supercritical birth and death process started from one individual at time zero. Each individual dies independently at rate $\mu$, and each individual gives birth to another individual at rate $\lambda$. If this population does not die out quickly, then in the long run it will grow exponentially at the rate $r:=\lambda-\mu$. Suppose that, after a large time $T$, we sample $n$ individuals uniformly at random from the population. We assume that, from the genetic data, we can recover the genealogical tree of these $n$ individuals. Our objective is to estimate the growth rate $r$ from this ancestral tree. This problem is of particular interest in the study of cancer, where the birth and death process models a growing population of cancer cells and one wants to estimate the rate at which a clone is expanding.
@@
Earlier work on estimating the growth rate used computationally intensive methods. [14] developed an approach based on Markov Chain Monte Carlo (MCMC) for estimating the growth rate of a birth and death process when each individual alive at time $T$ is sampled with some fixed probability. [16] and [10] used an approach called Phylofit to estimate the net growth rate. Phylofit is based on computing the joint likelihood of the coalescence times, for a model in which the population size levels off after an initial period of exponential growth.
@@
Recent theoretical developments have made it possible to estimate the growth rate by analytical methods. [2] computed the asymptotic site frequency spectrum of a sample of $n$ individuals from a birth and death process at a large time $T$. [4] and [8] developed a method to describe the exact genealogy of a sample of size $n$ from a birth and death process at time $T$. Using these results, [7] proposed two closely related methods of obtaining point and interval estimates for the growth rate $r$ without using MCMC. One method, which is entirely analytical, uses the lengths of the internal branches in the coalescent tree (the branches that are on the ancestral line of more than one lineage) to estimate the growth rate, while the other method involves maximum likelihood estimation.
@@
[7] demonstrate that their methods estimate the growth rate well for large sample sizes, providing comparable accuracy to previous methods ([14, 16, 10]) at much less computational cost. They apply their methods to blood cancer data, estimating the growth rate in 42 different clones and obtaining similar results to those obtained with more computationally intensive methods.
@@
Nevertheless, because the methods proposed by [7] are derived using asymptotics for large $n$, they do not perform as well as the more computationally intensive methods in small samples. Also, the estimator based on internal branch lengths depends not only on the coalescence times but also on the shape of the genealogical tree. This adds variability to the estimate because the tree shape contains no information about the growth rate $r$. Indeed, regardless of the growth rate, when one traces back the genealogy of $n$ sampled lineages in a birth and death process, at any time each remaining pair of lineages is equally likely to be the next pair that merges.
@@
In this paper, we propose a modification of the estimators in [7] which is not built on the assumption that $n$ is large. We avoid using information about the tree shape, which reduces the variance of the estimator. We compare the new approach to previous approaches using both simulated data and real data. We also show analytically that our estimator has, even for large $n$, less variability than the estimator based on internal branch lengths proposed by [7].
@@ 2.1 Internal and external branch lengths
Consider the genealogical tree of a sample of $n$ individuals from a birth and death process at time $T$. Let $L_{k,n}$ denote the sum of the lengths of all branches that are ancestral to $k$ of the $n$ sampled individuals. [2] showed that as $T\rightarrow\infty$, we have
@@
Branches that are ancestral to two or more sampled individuals are said to be internal, while branches that are ancestral to only one of the sampled individuals are said to be external (see Figure 1). Therefore, if we denote by $L^{in}$ the total length of the internal branches, then Durrett’s results imply that as $T\rightarrow\infty$, we have
@@
Note that this expression is inversely proportional to $r$, which implies that long internal branch lengths indicate a low growth rate, whereas short internal branch lengths indicate a higher growth rate. Based on this result, [7] proposed the estimator
@@
[7] also derived further theoretical properties of this estimator, which they used to obtain asymptotically valid confidence intervals for $r$. [12] then used similar methods to obtain second-order asymptotics for the full site frequency spectrum.
$H_{1}$$H_{2}$$H_{3}$$H_{4}$$H_{5}$$H_{6}$$H_{7}$$H_{8}$TT0
Figure 1: Genealogy of a sample of size $n=9$ from a birth and death process. The red lines represent internal branches. The blue lines represent external branches. The 8 coalescence times are indicated by $H_{1},\dots,H_{8}$.
@@ 2.2 Exact and approximate distributions of the coalescence times
For a genealogical tree of a sample of size $n$ with only pairwise mergers of ancestral lines, there are n1n-1 coalescence times. [4] and [8] showed that there is a relatively simple way to represent these coalescence times for the genealogy of a sample of size $n$ from a supercritical birth and death process at time $t$.
@@
We can then obtain the coalescence times $(H_{i,n,T})_{i=1}^{n-1}$ in the following way:
@@
Given these coalescence times, one can obtain the genealogical tree in two equivalent ways. One method, which is implicit in the work of [4], is to work backwards starting with $n$ lineages at time $T$ and merge two lineages, chosen uniformly at random, at each coalescence time. Another method, proposed by [8], is to use the coalescent point process, an approach which originated in [11] and was developed further by [1], [14], and [9]. We draw a vertical line of height $T$, followed by vertical lines of heights $H_{1,n,T},\dots,H_{n-1,n,T}$. At the top of each vertical line, we draw a horizontal dashed line to the left, stopping when we hit another vertical line. This procedure for generating a coalescent tree is illustrated in Figure 1.
[7] derived two successive approximations to these coalescence times, first taking a limit as $T\rightarrow\infty$ while $n$ is fixed, and then taking a limit as $n\rightarrow\infty$. The first of these limits yields the following way to approximate the coalescence times.
@@
Then taking the limit as $n\rightarrow\infty$ yields the following approach:
@@
    $f_{U_{i}}(u):=\frac{e^{u}}{(1+e^{u})^{2}},\quad u\in(0,\infty).$ (2.4)
@@
While the work of [7] focused on the second of these limits, we will consider primarily the first of these limits, in which $n$ is fixed and $T\rightarrow\infty$. This approach has the benefit that the resulting estimates work well for small sample sizes, while the distribution of $H_{i,n}$ remains mathematically tractable.
Based on (2.5), [7] also proposed a Maximum-Likelihood-Estimator (MLE) to estimate $r.$ They wrote $H_{i}=a+bU_{i}$ with $b=1/r$ and $a=T-\frac{1}{r}\left(\log(1/W)+\log(n)\right).$ Then they used the R-package maxLik ([5]) to calculate the MLE for $b.$ This estimator is called $\hat{r}_{MLE}.$
@@ 2.3 A modified estimator for the growth rate
We will consider here the second of the two approximations discussed in the previous section. Recall that the estimator $\hat{r}_{Lengths}$ in (2.1) was based on the internal branch length. Referring to Figure 1, the genealogical tree obtained using the coalescent point process has a branch on the left of height $T$, which we call the 0th branch, and branches of heights $H_{1,n},\dots,H_{n-1,n}$. For $i=1,\dots,n-2$, the length of the internal portion of the $i$th branch is $(H_{i,n}-H_{i+1,n})^{+}$, where $r^{+}=\max\{r,0\}$ for a real number $r$. The (n1)(n-1)st branch is entirely external, and the internal portion of the 0th branch is $\max_{1\leq i\leq n-1}H_{i,n}-H_{1,n}$. Therefore, the total internal branch length is
$L_{n}^{in}=\bigg{(}\max_{1\leq i\leq n-1}H_{i,n}-H_{1,n}\bigg{)}+\sum_{i=1}^{n-2}(H_{i,n}-H_{i+1,n})^{+}.$
Note, however, that $L^{in}$ depends on the order in which the coalescence times $H_{1,n},\dots,H_{n-1,n}$ are listed, even though, as can be seen from the construction in [4], the order of the coalescence times provides no information about the growth rate $r$. More formally, if we denote by $H_{n}^{(1)}\geq\dots\geq H_{n}^{(n-1)}$ the order statistics of the coalescence times, then the vector $(H_{n}^{(1)},\dots,H_{n}^{(n-1)})$ is a sufficient statistic for estimating $r$, as all (n1)!(n-1)! orders of the coalescence times are equally likely regardless of the value of $r$. Therefore, by the Rao-Blackwell Theorem, it should be possible to obtain a lower variance estimator by conditioning on the order statistics.
@@
$\mathbb{E}\left[L_{n}^{in}|H_{n}^{(1)},\ldots,H_{n}^{(n-1)}\right]=\bigg{(}H^{(1)}_{n}-\frac{1}{n-1}\sum_{i=1}^{n-1}H_{i,n}\bigg{)}+\frac{1}{n-1}\sum_{i=1}^{n-1}\sum_{j=1}^{n-1}(H_{i,n}-H_{j,n})^{+}.$ (2.6)
The second term on the right-hand side of (2.6) dominates when $n$ is large. Although the first term is relevant for small $n$ when considering the internal branch length, we have no reason to expect that including this term should lead to a better estimator for $r$. Therefore, we consider estimators of $r$ which are inversely proportional to the double sum in (2.6). That is, we consider estimators of the form
$\hat{r}=\frac{c(n)(n-1)(n-2)}{\sum_{i=1}^{n-1}\sum_{j=1}^{n-1}(H_{i,n}-H_{j,n})^{+}},$
where $c(n)$ is a constant depending only on $n$. There are (n1)(n2)(n-1)(n-2) terms in the denominator with $i\neq j$, and $\hat{r}$ equals $c(n)$ times the reciprocal of the average of these terms. The denominator can equivalently be written as $\sum_{i=1}^{n-1}\sum_{j=i+1}^{n-1}|H_{i,n}-H_{j,n}|$. Finally, note that when this estimator is applied to real data, the coalescence times derived from DNA sequence data will be used in place of $H_{i,n}$ and $H_{j,n}$.
@@ 2.4 The choice of $c(n)$
It was shown in [7] that $\mathbb{E}[(U_{i}-U_{j})^{+}]=1$, where $U_{i}$ and $U_{j}$ are independent with density given by (2.4). Therefore, we would expect $c_{MSE}(n)$, $c_{Bias}(n)$, and $c_{Inv}(n)$ all to tend to 11 as $n\rightarrow\infty$. The exact value of $c_{Inv}(n)$ can be calculated analytically. We will show in the Appendix that
@@ 2.5 Confidence Intervals
As in previous work ([14], [16], [10], [7]), we can obtain a confidence interval for the growth rate $r$ in addition to obtaining point estimates. Through simulation, we can find the $0.025$ and $0.975$ quantiles of the distribution of $S_{n}$, denoted respectively by $q_{0.025}$ and $q_{0.975}$, such that $\mathbb{P}(q_{0.025}<S_{n}<q_{0.975})=0.95$. Then, if we take $\hat{r}$ to be the estimator with $c(n)=1$, so that $\hat{r}=rS_{n}$, we get

### v2

@@ Abstract
The problem of estimating the growth rate of a birth and death processes based on the coalescence times of a sample of $n$ individuals has been considered by several authors ([stadler2009incomplete, williams2022life, mitchell2022clonal, Johnson2023]). This problem has applications, for example, to cancer research, when one is interested in determining the growth rate of a clone.
Recently, [Johnson2023] proposed an analytical method for estimating the growth rate using the theory of coalescent point processes, which has comparable accuracy to more computationally intensive methods when the sample size $n$ is large. We use a similar approach to obtain an estimate of the growth rate that is not based on the assumption that $n$ is large.
We demonstrate, through simulations using the R package cloneRate, that our estimator of the growth rate performs well in comparison with previous approaches when $n$ is small.
@@ 1 Introduction
Consider a birth and death process started from one individual at time zero. Each individual dies independently at rate $\mu$, and each individual gives birth to another individual at rate $\lambda$. If this population does not die out quickly, then in the long run it will grow exponentially at the rate $r:=\lambda-\mu$. We assume that $r>0$, which means the birth and death process is supercritical. Suppose that, after a large time $T$, we sample $n$ individuals uniformly at random from the population. Our objective is to estimate the growth rate $r$ from the genealogical tree of these $n$ sampled individuals. We assume that mutations are neutral, so that the birth and death rates are the same for all individuals.
In applications, there are various methods available for estimating the genealogical tree from whole-genome single-cell DNA data at a single time point (e.g. [de2013phylogenetic, kang2022sieve, kozlov2022cellphy, drummond2007beast, schliep2011phangorn, zafar2017sifit]). In the sequel, we assume that the genealogical tree, including the coalescence times when ancestral lines merge, can be recovered exactly. Note that if the birth, death, and mutation rates are multiplied by $c$, while $T$ is multiplied by $1/c$, the expected pattern of mutations observed in the sample would be unchanged. Therefore, reconstruction of the tree requires additional information that can be used to set the time axis. This could come from knowledge of the mutation rate or, in the case of clinical data such as that analyzed by [Johnson2023], the patient’s age.
The problem of estimating the growth rate is of particular interest in the study of cancer, where the birth and death process models a growing population of cancer cells and one wants to estimate the rate at which a clone is expanding. The growth rate is an important characteristic of cancer which can be interpreted as a measure of fitness ([brown2023updating]). Understanding the growth rate may help determine appropriate intervals for cancer screening ([peer1993age]), and may inform treatment decisions and contribute to improved cancer prevention strategies ([Johnson2023]). Beyond oncology, the concept of growth rate also plays an important role in epidemiology, where it reflects the speed at which a virus spreads through a population ([stadler2011inferring, stadler2013uncovering]).
@@
Earlier work on estimating the growth rate used computationally intensive methods. [stadler2009incomplete] developed an approach based on Markov Chain Monte Carlo (MCMC) for estimating the growth rate of a birth and death process when each individual alive at time $T$ is sampled with some fixed probability. [williams2022life] and [mitchell2022clonal] used an approach called Phylofit to estimate the growth rate. Phylofit is based on computing the joint likelihood of the coalescence times, i.e. the times when the ancestral lineages of the sampled cells merge, for a model in which the population size levels off after an initial period of exponential growth.
@@
Recent theoretical developments have made it possible to estimate the growth rate by analytical methods. [Durrett2013] computed the site frequency spectrum for a supercritical birth and death process, that is, the expected number of mutations appearing on $k$ out of $n$ individuals sampled at a large time $T$. [harris2020] and [LAMBERT201830] developed a method to describe the exact genealogy of a sample of size $n$ from a birth and death process at time $T$. Using these results, [Johnson2023] proposed two closely related methods of obtaining point and interval estimates for the growth rate $r$ without using MCMC. One method, which is entirely analytical, uses the lengths of the internal branches which can be calculated from the coalescence times in the coalescent tree to estimate the growth rate, while the other method involves maximum likelihood estimation.
@@
[Johnson2023] demonstrate that their methods estimate the growth rate well for large sample sizes, providing comparable accuracy to previous methods ([stadler2009incomplete, williams2022life, mitchell2022clonal]) at much less computational cost. They apply their methods to blood cancer data, estimating the growth rate in 42 different clones and obtaining similar results to those obtained with more computationally intensive methods.
@@
Nevertheless, because the methods proposed by [Johnson2023] are derived using asymptotics for large $n$, they do not perform as well as the more computationally intensive methods in small samples. Also, the estimator based on internal branch lengths depends not only on the coalescence times but also on the shape of the genealogical tree. This adds variability to the estimate because the tree shape contains no information about the growth rate $r$. Indeed, regardless of the growth rate, when one traces back the genealogy of $n$ sampled lineages in a birth and death process, at any time each remaining pair of lineages is equally likely to be the next pair that merges.
@@
In this paper, we propose a modification of the estimators in [Johnson2023] which is not built on the assumption that $n$ is large. We avoid using information about the tree shape, which reduces the variance of the estimator. We compare the new approach to previous approaches using both simulated data and real data. We also show analytically that our estimator has, even for large $n$, less variability than the estimator based on internal branch lengths proposed by [Johnson2023].
@@ 2.1 Internal and external branch lengths
Consider the genealogical tree of a sample of $n$ individuals from a birth and death process at time $T$. Let $L_{k,n}$ denote the sum of the lengths of all branches that are ancestral to $k$ of the $n$ sampled individuals. [Durrett2013] showed that as $T\rightarrow\infty$, we have
@@
Branches that are ancestral to two or more sampled individuals are said to be internal, while branches that are ancestral to only one of the sampled individuals are said to be external (see Figure 1 for a visualization). Therefore, if we denote by $L^{in}$ the total length of the internal branches, then Durrett’s results imply that as $T\rightarrow\infty$, we have
@@
Note that this expression is inversely proportional to $r$, which implies that long internal branch lengths indicate a low growth rate, whereas short internal branch lengths indicate a higher growth rate. Based on this result, [Johnson2023] proposed the estimator
@@
[Johnson2023] also derived further theoretical properties of this estimator, which they used to obtain asymptotically valid confidence intervals for $r$. [schweinsberg2025asymptotics] then used similar methods to obtain second-order asymptotics for the full site frequency spectrum.
[…]
Figure 1: Genealogy of a sample of size $n=9$ from a birth and death process. The green lines represent internal branches. The blue lines represent external branches. The 8 coalescence times are indicated by $H_{1},\dots,H_{8}$.
@@ 2.2 Exact and approximate distributions of the coalescence times
For a genealogical tree of a sample of size $n$ from a birth and death process, there are n1n-1 coalescence times. [harris2020] and [LAMBERT201830] showed that there is a relatively simple way to represent these coalescence times for the genealogy of a sample of size $n$ from a supercritical birth and death process at time $T$.
@@
We can then generate the distribution of the coalescence times $(H_{i,n,T})_{i=1}^{n-1}$ in the following way:
@@
Given these coalescence times, one can generate the genealogical tree in two equivalent ways. One method, which is implicit in the work of [harris2020], is to work backwards starting with $n$ lineages at time $T$ and merge two lineages, chosen uniformly at random, at each coalescence time. Another method, proposed by [LAMBERT201830], is to use the coalescent point process, an approach which originated in [popovic2004] and was developed further by [Aldous_Popovic_2005], [stadler2009incomplete], and [lambert2013birth]. We draw a vertical line of height $T$, followed by vertical lines of heights $H_{1,n,T},\dots,H_{n-1,n,T}$. At the top of each vertical line, we draw a horizontal dashed line to the left, stopping when we hit another vertical line. This procedure for generating a coalescent tree is illustrated in Figure 1.
[Johnson2023] derived two successive approximations to these coalescence times, first taking a limit as $T\rightarrow\infty$ while $n$ is fixed, and then taking a limit as $n\rightarrow\infty$. The first of these limits yields the following way to approximate the coalescence times.
@@
After taking the limit as $T\rightarrow\infty$, we can take the limit as $n\rightarrow\infty$. Then the distribution of the coalescent times converges to a distribution that can be represented in the following way:
@@
    $f_{U_{i}}(u):=\frac{e^{u}}{(1+e^{u})^{2}},\quad u\in(-\infty,\infty).$ (2.4)
@@
While the work of [Johnson2023] focused on the second of these limits, we will consider primarily the first of these limits, in which $n$ is fixed and $T\rightarrow\infty$, i.e. $(H_{i,n})_{i=1,...,n-1}$. This approach has the benefit that the resulting estimates work well for small sample sizes, while the distribution of $H_{i,n}$ remains mathematically tractable.
Based on (2.5), [Johnson2023] also proposed a maximum likelihood estimator (MLE) to estimate $r.$ They wrote $H_{i}=a+bU_{i}$ with $b=1/r$ and $a=T-\frac{1}{r}\left(\log(1/W)+\log(n)\right).$ Since $U_{i}$ is symmetric, this is equivalent to setting $H_{i}=a-bU_{i}.$ Then they used the R-package maxLik ([henningsen2011maxlik]) to calculate the MLE for $b.$ The reciprocal of this estimator is denoted by $\hat{r}_{MLE}.$
@@ 2.3 A modified estimator for the growth rate
We will consider here the second of the two approximations discussed in the previous section. Recall that the estimator $\hat{r}_{Lengths}$ in (2.1) was based on the internal branch length. Referring to Figure 1, the genealogical tree obtained using the coalescent point process has a branch on the left of height $T$, which we call the 0th branch, and branches of heights $H_{1,n},\dots,H_{n-1,n}$. For $i=1,\dots,n-2$, the length of the internal portion of the $i$th branch is $(H_{i,n}-H_{i+1,n})^{+}$, where $x^{+}=\max\{x,0\}$ for a real number $x$. The (n1)(n-1)st branch is entirely external, and the internal portion of the 0th branch is $\max_{1\leq i\leq n-1}H_{i,n}-H_{1,n}$. Therefore, the total internal branch length is
$L_{n}^{in}=\bigg(\max_{1\leq i\leq n-1}H_{i,n}-H_{1,n}\bigg)+\sum_{i=1}^{n-2}(H_{i,n}-H_{i+1,n})^{+}.$ (2.6)
Note, however, that $L^{in}$ depends on the order in which the coalescence times $H_{1,n},\dots,H_{n-1,n}$ are listed, even though, as can be seen from the construction in [harris2020], the order of the coalescence times provides no information about the growth rate $r$. More formally, if we denote by $H_{n}^{(1)}\leq\dots\leq H_{n}^{(n-1)}$ the order statistics of the coalescence times, then the vector $(H_{n}^{(1)},\dots,H_{n}^{(n-1)})$ is a sufficient statistic for estimating $r$, as all (n1)!(n-1)! orders of the coalescence times are equally likely regardless of the value of $r$. Therefore, by the Rao-Blackwell Theorem, it should be possible to obtain an estimator with lower or equal variance by conditioning on the order statistics.
@@
$\mathbb{E}\left[L_{n}^{in}|H_{n}^{(1)},\ldots,H_{n}^{(n-1)}\right]$ $=\bigg(H^{(n-1)}_{n}-\frac{1}{n-1}\sum_{i=1}^{n-1}H_{i,n}\bigg)+\frac{1}{n-1}\sum_{i=1}^{n-1}\sum_{j=1}^{n-1}(H_{i,n}-H_{j,n})^{+}$ (2.7)
$=\bigg(H^{(n-1)}_{n}-\frac{1}{n-1}\sum_{i=1}^{n-1}H_{i,n}\bigg)+\frac{1}{n-1}\sum_{i=1}^{n-1}\sum_{j=i+1}^{n-1}|H_{i,n}-H_{j,n}|.$
A detailed derivation of (2.7) can be found in Appendix 5.1. The second term on the right-hand side of (2.7) dominates when $n$ is large, which suggests that one can obtain a good estimator of $r$ that is based on the average of the terms $(H_{i,n}-H_{j,n})^{+}$. Note that including the first term would double the weight assigned to the terms $(H_{i,n}-H_{j,n})^{+}$ in which $H_{i,n}$ is the largest of $H_{1,n},\dots,H_{n-1,n}$, but we prefer to pursue the simpler approach of treating all of the terms $H_{i,n}-H_{j,n}$ symmetrically. Therefore, we consider estimators of $r$ which are inversely proportional to the double sum in (2.7). That is, we consider estimators of the form
$\hat{r}=\frac{c(n)(n-1)(n-2)}{\sum_{i=1}^{n-1}\sum_{j=1}^{n-1}(H_{i,n}-H_{j,n})^{+}},$ (2.8)
where $c(n)$ is a constant depending only on $n$. There are (n1)(n2)(n-1)(n-2) terms in the denominator with $i\neq j$, and $\hat{r}$ equals $c(n)$ times the reciprocal of the average of these terms. Note that when this estimator is applied to real data, the coalescence times derived from DNA sequence data will be used in place of $H_{i,n}$ and $H_{j,n}$.
@@ 2.4 The choice of $c(n)$
It was shown in [Johnson2023] that $\mathbb{E}[(U_{i}-U_{j})^{+}]=1$, where $U_{i}$ and $U_{j}$ are independent with density given by (2.4). Therefore, we would expect $c_{MSE}(n)$, $c_{Bias}(n)$, and $c_{Inv}(n)$ all to tend to $1$ as $n\rightarrow\infty$. The exact value of $c_{Inv}(n)$ can be calculated analytically. We will show in the Appendix that
@@ 2.5 Confidence Intervals
As in previous work ([stadler2009incomplete], [williams2022life], [mitchell2022clonal], [Johnson2023]), we can obtain a confidence interval for the growth rate $r$ in addition to obtaining point estimates. Through simulation, we can find the $0.025$ and $0.975$ quantiles of the distribution of $S_{n}$, denoted respectively by $q_{0.025}$ and $q_{0.975}$, such that $\mathbb{P}(q_{0.025}<S_{n}<q_{0.975})=0.95$. Then, if we take $\hat{r}$ to be the estimator with $c(n)=1$, so that $\hat{r}=rS_{n}$, we get